The data show the number of specific recorded cases of meningitis for 14 specific states.

10, 1, 1, 28, 15, 41, 4, 4, 8, 2, 3, 53, 34, 1

Find the z values for each

a.10

b.28

c.41
Answer #1

From the given data,

Mean = 14.6429

Standard deviation = 17.2431

z-score = X - Mean / SD

a)

z = 10 - 14.6429 / 17.2431

= -0.27

b)

z = 28 - 14.6429 / 17.2431

= 0.77

c)

z = 41 - 14.6429 / 17.2431

= 1.53